Constants Part A If a curve with a radius of 87 m is properly banked for a car traveling 79 km/h, what must be the coefficient of static friction for a car not to skid when traveling at 103 km/h? Express your answer using two significant figures. Hanin-0.1627 Submit Previous Answers Request Answer Incorrect; Try Again; 3 attempts remaining | Provide Feedback Next >
Answer #1

A] Here v = 79 km/h = 79*5/18 m/s

tan theta = v^2/rg = [79*5/18]^2/[87*9.8] = 0.5648

speed with friction = sqrt(rg* (tan theta + u)/(1 - u tan theta))

103*5/18 = sqrt((87*9.8*(0.5648+u)/(1-u*0.5648))

u = 0.26 answer